# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

### v1

@@ III.2.2 Results (symmetric setup, identical contacts)
$\mathcal{F}^{\text{CI}}(\delta\phi,\gamma)$ $\approx\frac{2-x^{2}+x^{4}}{2(1+x^{2})^{2}}\,,$
Hence, the value of $\mathcal{F}^{\text{CI}}$ depends on the order of taking the limits $\delta\phi\to 0$ and $\gamma\to 0,$ so that if we fix $\delta\phi$ and tend $\gamma$ to zero, then instead of Eq. (23) we get the usual expression for the tunnel contact:
$\mathcal{F}^{\text{CI}}\to 1/2,\quad\gamma\to 0,\quad\delta\phi~{}\mbox{ is fixed}.$
From Eqs. (24) we can get an analytical expression describing the orange line in Fig. 7a. To this end, we fix $\delta\phi$ and change tunneling coupling from $\gamma=0$ to relatively large value: $\delta\phi\ll\gamma\ll 1.$ Then, we get
$\mathcal{F}^{\text{CI}}\approx\frac{3}{4}-\frac{1}{2}\frac{\mathcal{T}^{\text{CI}}}{\pi\delta\phi}\pm\frac{1}{4}\sqrt{1-\left(\frac{\mathcal{T}^{\text{CI}}}{\pi\delta\phi}\right)^{2}},\quad 0<\mathcal{T}^{\text{CI}}<\pi|\delta\phi|,$ (25)
The sign in front of the last term distinguish two branches of the orange curve. The minimum of $\mathcal{T}^{\rm CI}$ is reached at the lower branch at $\mathcal{T}^{\rm CI}=15\pi|\delta\phi|/16$ and is given by $\mathcal{F}^{\rm CI}=17/32,$ , which is the value of the dashed line in Fig. 7a for $\mathcal{T}^{\text{CI}}=0$.
In Fig. 7a we fixed $\phi$ and changed $\gamma$ from zero to infinity. Fixing $\gamma$ instead and changing $\phi$ in the interval $10<\phi<$ 1, we find the following expression connecting transmission coefficient and the Fano factor valid for any $\gamma$ in the interval $0<\gamma<\infty$:
$\mathcal{F}^{\text{CI}}=1-\frac{5+\gamma^{2}}{2}~{}\frac{\mathcal{T}^{\text{CI}}}{2\gamma}+2\left(\frac{\mathcal{T}^{\text{CI}}}{2\gamma}\right)^{2},$ (26)
$0<\mathcal{T}^{\text{CI}}<\frac{2\gamma}{1+\gamma^{2}}.$
This equation has minimum at $\mathcal{T}^{\text{CI}}_{\rm min}=(5+\gamma^{2})\gamma/4$ given by $\mathcal{F}^{\text{CI}}_{\rm min}=[7-\gamma^{2}(10+\gamma^{2})]/32.$ The family of the curves described by Eq. (26) for different $\gamma$ is shown in Fig. 7b. Note that the dashed curves in panels (a) and (b) of Fig. 7 coincide.
@@ III.2.3 Results (asymmetric setup)
The asymmetry of setup becomes important when $a\gtrsim k_{F}^{-1}.$ For not too large voltages, $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in scattering amplitudes. The expression for the transmission coefficient in asymmetric setup was obtained earlier in Ref. Dmitriev et al. (2010). Analytical expression for the Fano factor in this case can be found by using formulas presented in Appendixes A and B. Its general form is too cumbersome and we present it here only for the flux values close to integer and half-integer number and for weak tunneling coupling.
@@
$\mathcal{F}^{\text{CI}}_{1/2}(\delta\phi,\gamma)=\frac{2-\bar{x}^{2}+\bar{x}^{4}}{2(1+\bar{x}^{2})^{2}}+\frac{\sin^{2}(k_{F}a)[3\bar{x}^{2}-\sin^{2}(k_{F}a)]}{2(1+\bar{x}^{2})[\bar{x}^{2}+\sin^{2}(k_{F}a)]},$ (27)
@@
Similar to Eq. (24), the value of $\mathcal{F}^{\text{CI}}_{1/2}$ at $\delta\phi\to 0$, $\gamma\to 0$ depends on the order of limits. Fixing $\delta\phi$ and decreasing the tunneling coupling to zero, $\gamma\to 0,$ we get $\bar{x}\to\infty,$ so that $\mathcal{F}^{\text{CI}}_{1/2}\to 1/2$. Putting $\delta\phi$ to zero first, we find $\mathcal{F}^{\text{CI}}_{1/2}\to 1-(1/2)\sin^{2}(k_{\rm F}a).$
@@
The dependencies of ${\mathcal{T}}^{\text{CI}}$ and $\mathcal{F}^{\text{CI}}$ on the magnetic flux, $\phi$, are shown in Fig. 6c and 6d. Comparing these figures with Fig. 6a and 6b plotted for symmetric case, we see that new peaks at integer values of the flux appear for $a\neq 0$ both in the transmission coefficient and in the Fano factor.
@@ III.2.4 Results (different tunneling couplings)
In this subsection, we present results for the case of symmetric setup, $a=0,$ but different right and left contacts, characterized by parameter, $\gamma_{\text{R}}$ and $\gamma_{\text{L}}$, respectively. Using general equations presented in Appendix B and performing standard energy averaging with the use of formulas from Appendix A, one can find the following expressions for the transmission coefficient and Fano factor:
${\mathcal{T}}^{\text{CI}}(\phi,\gamma_{\text{R}},\gamma_{\text{L}})=\kappa^{-1/2}\frac{2\bar{\gamma}\cos^{2}(\pi\phi)}{\bar{\gamma}^{2}+\cos^{2}(\pi\phi)}\,,$ (28)
$\mathcal{F}^{\text{CI}}(\phi,\gamma_{\text{R}},\gamma_{\text{L}})=1-\kappa^{-1}\frac{\cos^{2}(\pi\phi)}{2(\bar{\gamma}^{2}+\cos^{2}(\pi\phi))^{2}}$
×$\times\left((1+4\kappa)\bar{\gamma}^{2}+\bar{\gamma}^{4}+(1+\bar{\gamma}^{2})\cos^{2}(\pi\phi)\right)\,,$
@@
$\bar{\gamma}=\sqrt{\gamma_{\text{L}}\gamma_{\text{R}}}\,,\quad\kappa=(\gamma_{\text{L}}+\gamma_{\text{R}})^{2}/4\gamma_{\text{L}}\gamma_{\text{R}}\,.$
@@
$\mathcal{F}^{\text{CI}}=1-\frac{1}{2\kappa}-\frac{\kappa}{2}\left({\mathcal{T}}^{\text{CI}}\right)^{2}\,.$ (29)
@@ III.3 Helical AB interferometer
Next, we discuss the Fano factor of the current noise in HI having two arms with lengths $L_{1}$ and $L_{2}.$ Just as in the case of CI we start formula for energy dependent amplitude $\hat{t}^{\rm HI}(\varepsilon),$ which is a matrix in spinful case, and then perform energy averaging. The analytical expression obtained for $\mathcal{F}^{\rm HI}$ is rather cumbersome, so we present it in the Appendix C focusing in the main text on the most interesting limiting cases.
@@ III.3.1 Scattering matrix
$\begin{pmatrix}a^{\uparrow}\\ a^{\downarrow}\end{pmatrix}=\hat{t}^{\rm HI}\begin{pmatrix}b^{\uparrow}\\ b^{\downarrow}\end{pmatrix}\,,$ (30)
where $(b^{\uparrow},b^{\downarrow})$ and $(a^{\uparrow},a^{\downarrow})$ are the amplitudes of incoming (from the left contact) and outgoing (to the right contact) waves, respectively (see Fig. 1). This matrix has been obtained earlier Niyazov et al. (2020):
$\hat{t}^{\rm CI}=$ $\frac{r^{2}e^{2\pi i\phi L_{1}/L}}{t^{2}}\begin{pmatrix}e^{ikL_{1}}&0\\ 0&e^{-ikL_{1}}\end{pmatrix}\begin{pmatrix}t&0\\ 0&1\end{pmatrix}\hat{g}\begin{pmatrix}t&0\\ 0&1\end{pmatrix}\,,$ (31)
$\hat{g}=$ $\cos\theta\left[\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}+\frac{1}{2}\sum\limits_{\alpha=\pm 1}\frac{1+\alpha\hat{H}}{1-t^{2}e^{i(kL+\alpha 2\pi\phi_{0})}}\right]\,.$
@@
$\hat{H}=\begin{pmatrix}a&be^{i\xi}\\ be^{-i\xi}&-a\end{pmatrix}\,,$ (33)
@@
Calculations shows that asymmetry of the arms does not play any role in the HI. Also, position of MD $x_{0}$ drops out from final expressions for $\mathcal{T}^{\rm HI}$ and $\mathcal{F}^{\rm HI}.$ We therefore consider two cases: identical contacts (see next subsection and Appendix C.1) and non-identical contacts (see Appendix C.2)
@@
where the parameter describing the difference of tunneling contacts [see also App. C.2] is
$\beta=\tfrac{1}{2}\ln(t_{\text{R}}/t_{\text{L}})\,.$ (36)
Figure 8: Panel (a): The dependence of the Fano factor, $\mathcal{F}^{\text{HI}}$, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$; for $r\to 1$ there an is optimal value, $R_{\theta}=R_{\theta}^{\rm opt},$ where the Fano factor has maximum. Panel (b): Cross-sections of panel (a) at different values of $R_{\theta};$ dependence on the flux is weak and is illustrated for $R_{\theta}=R_{\theta}^{\rm opt}$ by broadening of the curve (shown by blue color) for flux belonging to interval $10<\phi<$ 1.
@@ III.3.2 Results (identical contacts)
The transmission coefficient and the Fano factor are obtained with the use of Eqs. (9), (13),(14), (15), and (31). Direct energy averaging in Eqs. (13) and (14) yields
${\mathcal{T}}^{\text{HI}}=\tanh\lambda\left[\!1-\frac{\sin^{2}\theta\sinh^{2}\lambda\cosh(2\lambda)}{\cosh^{2}(2\lambda)-\cos^{2}\theta\cos^{2}(2\pi\phi)}\!\right]$ (37)
and
${\mathcal{T}}_{2}^{\text{HI}}=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,.$ (38)
@@
and the coefficients $\tilde{A}_{(m,n)}$ depend on $R_{\theta},R_{\phi}$ do not depend on $\lambda.$ Analytical expressions for these coefficients are presented in the Appendix C. The Fano factor is connected with $\mathcal{T}^{\rm HI}$ and $\mathcal{T}^{\rm HI}_{2}$ by Eq. (15).
In Fig. 8a, we present dependence of the Fano factor, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$ for $\phi=0$. For “metallic” contact, $r\to 1,$ dependence on the backscattering strength $R_{\theta}$ is non-monotonous. Comparing this figure with analogous Fig. LABEL:FanoCI for CI, we see that in both cases the Fano factor has extremum as a function of squared backscattering amplitude. However, for CI the Fano factor has minimum, while for HI it has maximum.
@@
$\mathcal{F}_{\rm n}^{\text{HI}}(\phi)=\frac{\mathcal{F}^{\text{HI}}(\phi)-\mathcal{F}^{\text{HI}}(1/4)}{\mathcal{F}^{\text{HI}}(0)-\mathcal{F}^{\text{HI}}(1/4)}.$ (40)
The dependence of the normalized quantity $\mathcal{F}_{\rm n}^{\text{HI}}$ on the magnetic flux for several sets of parameters $R_{\theta}$ and $\lambda$ is shown in Fig. 9 along with the flux dependence of the normalized transmission coefficient. One can see sharp resonances both in $\mathcal{F}_{\rm n}^{\text{HI}}$ and $\mathcal{T}_{\rm n}^{\text{HI}}$ at weak tunneling coupling and small $R_{\theta}$ (see Fig. 9c) which evolve into harmonic Aharonov-Bohm oscillations with increasing $\lambda$ and $R_{\theta}$ (see Fig. 9d)
Figure 9: Dependence of the normalized factor Fano, $\mathcal{F}_{n}^{\text{HI}}$, and the normalized conductance, $\mathcal{T}_{n}^{\text{HI}}$, for the helical interferometer on magnetic flux, $\phi$, at different strengths of scattering by a magnetic defect $R_{\theta}$ and tunneling amplitude, $r$. (a) $r=0.7,R_{\theta}=1/8$ (b) $r=0.7,R_{\theta}=3/8$ (c) $r=0.4,R_{\theta}=1/1000$ (d) $r=0.7,R_{\theta}=1/2$.

### v2

@@ III.2.2 Results (symmetric setup, identical contacts)
$\mathcal{F}^{\mathrm{CI}}(\delta\phi,\gamma)$ $\approx\frac{2-x^{2}+x^{4}}{2(1+x^{2})^{2}}\,$, (27)
Hence, the value of $\mathcal{F}^{\mathrm{CI}}$ depends on the order of taking the limits $\delta\phi\to 0$ and $\gamma\to 0,$ so that if we fix $\delta\phi$ and tend $\gamma$ to zero, then instead of Eq. (25) we get the usual expression for the tunnel contact:
$\mathcal{F}^{\mathrm{CI}}\to 1/2,\quad\gamma\to 0,\quad\delta\phi~\mbox{ is fixed}.$
From Eqs. (27) we can get an analytical expression describing the orange line in Fig. 7a. To this end, we fix $\delta\phi$ and change tunneling coupling from $\gamma=0$ to relatively large value: $\delta\phi\ll\gamma\ll 1.$ Then, we get at $0<\mathcal{T}^{\mathrm{CI}}<\pi|\delta\phi|$:
$\mathcal{F}^{\mathrm{CI}}\approx\frac{3}{4}-\frac{1}{2}\frac{\mathcal{T}^{\mathrm{CI}}}{\pi\delta\phi}\pm\frac{1}{4}\sqrt{1-\left(\frac{\mathcal{T}^{\mathrm{CI}}}{\pi\delta\phi}\right)^{2}}\,.$ (28)
The sign in front of the last term distinguish two branches of the orange curve. The minimum of $\mathcal{T}^{\rm CI}$ is reached at the lower branch at $\mathcal{T}^{\rm CI}=15\pi|\delta\phi|/16$ and is given by $\mathcal{F}^{\rm CI}=17/32,$ , which is the value of the dashed line in Fig. 7a for $\mathcal{T}^{\mathrm{CI}}=0$.
In Fig. 7a we fixed $\phi$ and changed $\gamma$ from zero to infinity. Fixing $\gamma$ instead and changing $\phi$ in the interval $0<\phi<1$, we find the following expression connecting transmission coefficient and the Fano factor valid for any $\gamma$ in the interval $0<\gamma<\infty$:
$\mathcal{F}^{\mathrm{CI}}$ $=$ $1-\frac{5+\gamma^{2}}{2}~\frac{\mathcal{T}^{\mathrm{CI}}}{2\gamma}+2\left(\frac{\mathcal{T}^{\mathrm{CI}}}{2\gamma}\right)^{2},$
$0<\mathcal{T}^{\mathrm{CI}}<\frac{2\gamma}{1+\gamma^{2}}.$
This equation has minimum at $\mathcal{T}^{\mathrm{CI}}_{\rm min}=(5+\gamma^{2})\gamma/4$ given by $\mathcal{F}^{\mathrm{CI}}_{\rm min}=[7-\gamma^{2}(10+\gamma^{2})]/32.$ The family of the curves described by Eq. (III.2.2) for different $\gamma$ is shown in Fig. 7b. Note that the dashed curves in panels (a) and (b) of Fig. 7 coincide. From above discussion, we see that behavior of the noise of CI is very singular for $\phi$ close to $1/2.$ This means that noise is very sensitive to the flux in the vicinity of this point.
@@ III.2.3 Results (asymmetric setup)
The asymmetry of setup becomes important when $a\geq k_{F}^{-1}.$ For not too large voltages, $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in scattering amplitudes. The expression for the transmission coefficient in asymmetric setup was obtained earlier in Ref. Dmitriev et al. (2010). Analytical expression for the Fano factor in this case can be found by using formulas presented in A and B. Its general form is too cumbersome and we present it here only for the flux values close to integer and half-integer number and for weak tunneling coupling.
@@
$\mathcal{F}^{\mathrm{CI}}_{1/2}(\delta\phi,\gamma)$ $=\frac{2-\bar{x}^{2}+\bar{x}^{4}}{2(1+\bar{x}^{2})^{2}}$ (30)
$+\frac{\sin^{2}(k_{F}a)[3\bar{x}^{2}-\sin^{2}(k_{F}a)]}{2(1+\bar{x}^{2})[\bar{x}^{2}+\sin^{2}(k_{F}a)]},$
@@
Similar to Eq. (27), the value of $\mathcal{F}^{\mathrm{CI}}_{1/2}$ at $\delta\phi\to 0$, $\gamma\to 0$ depends on the order of limits. Fixing $\delta\phi$ and decreasing the tunneling coupling to zero, $\gamma\to 0,$ we get $\bar{x}\to\infty,$ so that $\mathcal{F}^{\mathrm{CI}}_{1/2}\to 1/2$. Putting $\delta\phi$ to zero first, we find
$\mathcal{F}^{\mathrm{CI}}_{1/2}\to 1-\frac{\sin^{2}(k_{\rm F}a)}{2}.$ (31)
@@
As seen from Eqs. (30) and (31), the asymmetry of the device strongly affect the noise intensity even for fixed magnetic flux, so that changing $a$ within a small interval about the Fermi wavelength, we change $\mathcal{F}^{\rm CI}$ by a factor on the order of two.
The dependencies of ${\mathcal{T}}^{\mathrm{CI}}$ and $\mathcal{F}^{\mathrm{CI}}$ on the magnetic flux, $\phi$, are shown in Fig. 6c and 6d. Comparing these figures with Fig. 6a and 6b plotted for symmetric case, we see that new peaks at integer values of the flux appear for $a\neq 0$ both in the transmission coefficient (antiresonances) and in the Fano factor (resonances) for integer values of flux. With increasing asymmetry, the amplitudes of the peaks corresponding to half-integer values of the flux decrease, while the amplitudes of the peaks at integer flux values increase. We conclude that asymmetry of the device, which is always present in real experiments Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007), strongly modifies shot noise.
@@ III.2.4 Results (different tunneling couplings)
In this subsection, we present results for the case of symmetric setup, $a=0,$ but different right and left contacts, characterized by parameter, $\gamma_{\mathrm{R}}$ and $\gamma_{\mathrm{L}}$, respectively. Using general equations presented in B and performing standard energy averaging with the use of formulas from A, one can find the following expressions for the transmission coefficient and Fano factor:
${\mathcal{T}}^{\mathrm{CI}}(\phi,\gamma_{\mathrm{R}},\gamma_{\mathrm{L}})=\kappa^{-1/2}\frac{2\bar{\gamma}\cos^{2}(\pi\phi)}{\bar{\gamma}^{2}+\cos^{2}(\pi\phi)}\,$,
$\mathcal{F}^{\mathrm{CI}}(\phi,\gamma_{\mathrm{R}},\gamma_{\mathrm{L}})=1-\kappa^{-1}\frac{\cos^{2}(\pi\phi)}{2(\bar{\gamma}^{2}+\cos^{2}(\pi\phi))^{2}}$
$\times\left((1+4\kappa)\bar{\gamma}^{2}+\bar{\gamma}^{4}+(1+\bar{\gamma}^{2})\cos^{2}(\pi\phi)\right)\,$,
@@
$\bar{\gamma}=\sqrt{\gamma_{\mathrm{L}}\gamma_{\mathrm{R}}}\,,\quad\kappa=(\gamma_{\mathrm{L}}+\gamma_{\mathrm{R}})^{2}/4\gamma_{\mathrm{L}}\gamma_{\mathrm{R}}\,.$
@@
$\mathcal{F}^{\mathrm{CI}}=1-\frac{1}{2\kappa}-\frac{\kappa}{2}\left({\mathcal{T}}^{\mathrm{CI}}\right)^{2}\,.$ (32)
@@ III.3 Helical AB interferometer
Next, we discuss the Fano factor of the current noise in HI having two arms with lengths $L_{1}$ and $L_{2}.$ Just as in the case of CI we start formula for energy dependent amplitude $\hat{t}^{\rm HI}(\varepsilon),$ which is a matrix in spinful case, and then perform energy averaging. The analytical expression obtained for $\mathcal{F}^{\rm HI}$ is rather cumbersome, so we present it in the C focusing in the main text on the most interesting limiting cases.
@@ III.3.1 Scattering matrix
$\left(\begin{array}[]{c}a^{\uparrow}\\ a^{\downarrow}\end{array}\right)=\hat{t}^{\rm HI}\left(\begin{array}[]{c}b^{\uparrow}\\ b^{\downarrow}\end{array}\right)\,$, (33)
where $(b^{\uparrow},b^{\downarrow})$ and $(a^{\uparrow},a^{\downarrow})$ are the amplitudes of incoming (from the right contact) and outgoing (to the left contact) waves, respectively (see Fig. 1). This matrix has been obtained earlier Niyazov et al. (2020):
$\hat{t}^{\rm HI}$ $=\frac{r^{2}e^{2\pi i\phi L_{1}/L}}{t^{2}}\left(\begin{array}[]{cc}e^{ikL_{1}}&0\\ 0&e^{-ikL_{1}}\end{array}\right)$ (34)
$\times\left(\begin{array}[]{cc}t&0\\ 0&1\end{array}\right)\hat{g}\left(\begin{array}[]{cc}t&0\\ 0&1\end{array}\right)\,$,
$\hat{g}$ $=\cos\theta\left[\left(\begin{array}[]{cc}0&0\\ 0&-1\end{array}\right)\right$.
$\left.+\frac{1}{2}\sum\limits_{\alpha=\pm 1}\frac{1+\alpha\hat{H}}{1-t^{2}e^{i(kL+\alpha 2\pi\phi_{0})}}\right]\,.$
@@
$\hat{H}=\left(\begin{array}[]{cc}a&be^{i\xi}\\ be^{-i\xi}&-a\end{array}\right)\,$, (36)
@@
Calculations shows that asymmetry of the arms does not play any role in the HI. Also, position of MD $x_{0}$ drops out from final expressions for $\mathcal{T}^{\rm HI}$ and $\mathcal{F}^{\rm HI}.$ We therefore consider two cases: identical contacts (see next subsection and C.1) and non-identical contacts (see C.2)
@@
where the parameter describing the difference of tunneling contacts [see also C.2] is
$\beta=\frac{1}{2}\ln(t_{\mathrm{R}}/t_{\mathrm{L}})\,.$ (39)
Figure 8: Panel (a): The dependence of the Fano factor, $\mathcal{F}^{\mathrm{HI}}$, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$; for $r\to 1$ there an is optimal value, $R_{\theta}=R_{\theta}^{\rm opt},$ where the Fano factor has maximum. Panel (b): Cross-sections of panel (a) at different values of $R_{\theta};$ dependence on the flux is weak and is illustrated for $R_{\theta}=R_{\theta}^{\rm opt}$ by broadening of the curve (shown by blue color) for flux belonging to interval $0<\phi<1$. (these figures were presented in Ref. Niyazov et al. (2024) in a slightly different way). The point $r=1$ is “noiseless” both for ballistic case $R_{\theta}=0$ and the case of very strong defect $R_{\theta}=1.$
@@ III.3.2 Results (identical contacts)
The transmission coefficient and the Fano factor are obtained with the use of Eqs. (10), (14), (15), (16), and (34). Direct energy averaging in Eqs. (14) and (15) yields Niyazov et al. (2021a)
${\mathcal{T}}^{\mathrm{HI}}$ $=\tanh\lambda\left[1\right$. (40)
$\left.-\frac{\sin^{2}\theta\sinh^{2}\lambda\cosh(2\lambda)}{\cosh^{2}(2\lambda)-\cos^{2}\theta\cos^{2}(2\pi\phi)}\right]$
and Niyazov et al. (2024)
${\mathcal{T}}_{2}^{\mathrm{HI}}$ $=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,.$ (41)
@@
and the coefficients $\tilde{A}_{(m,n)}$ depend on $R_{\theta},R_{\phi}$ do not depend on $\lambda.$ Analytical expressions for these coefficients are presented in the C.1. The Fano factor is connected with $\mathcal{T}^{\rm HI}$ and $\mathcal{T}^{\rm HI}_{2}$ by Eq. (16).
In Fig. 8a, we present dependence of the Fano factor, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$ for $\phi=0$. For “metallic” contact, $r\to 1,$ dependence on the backscattering strength $R_{\theta}$ is non-monotonous. Comparing this figure with analogous Fig. 5 for CI, we see that in both cases the Fano factor has extremum as a function of squared backscattering amplitude. However, for CI the Fano factor has minimum, while for HI it has maximum.
@@
$\mathcal{F}_{\rm n}^{\mathrm{HI}}(\phi)=\frac{\mathcal{F}^{\mathrm{HI}}(\phi)-\mathcal{F}^{\mathrm{HI}}(1/4)}{\mathcal{F}^{\mathrm{HI}}(0)-\mathcal{F}^{\mathrm{HI}}(1/4)}.$ (43)
The dependence of the normalized quantity $\mathcal{F}_{\rm n}^{\mathrm{HI}}$ on the magnetic flux for several sets of parameters $R_{\theta}$ and $\lambda$ is shown in Fig. 9 along with the flux dependence of the normalized transmission coefficient. One can see sharp resonances both in $\mathcal{F}_{\rm n}^{\mathrm{HI}}$ and $\mathcal{T}_{\rm n}^{\mathrm{HI}}$ at weak tunneling coupling and small $R_{\theta}$ (see Fig. 9c) which evolve into harmonic Aharonov-Bohm oscillations with increasing $\lambda$ and $R_{\theta}$ (see Fig. 9d)
Figure 9: Dependence of the normalized factor Fano, $\mathcal{F}_{n}^{\mathrm{HI}}$, and the normalized conductance, $\mathcal{T}_{n}^{\mathrm{HI}}$, for the helical interferometer on magnetic flux, $\phi$, at different strengths of scattering by a magnetic defect $R_{\theta}$ and tunneling amplitude, $r$. (a) $r=0.7,R_{\theta}=1/8$ (b) $r=0.7,R_{\theta}=3/8$ (c) $r=0.4,R_{\theta}=1/1000$ (d) $r=0.7,R_{\theta}=1/2$. Shape of the curve $\mathcal{F}^{\rm HI}(\phi)$ strongly depends on the transparency of the contacts.
@@
Fig. 10 is one of the key results concerning HI. Indeed, measuring simultaneously both conductance and noise and using Fig. 10), one can find the strength of the backscattering defect. It is worth stressing that measuring the conductance only is insufficient to find the strength of the defect, because the conductance depends also on the properties of the contacts, which are actually not known to a sufficient precision. Importantly, by using both conductance and noise measurements, together with Fig. 10 we can find $R_{\theta}$ without having detailed information about amplitudes $t$ and $r,$ characterizing the contacts. It can help to identify backscattering mechanism that apparently always exists in the realistic experimental structures Olshanetsky et al. (2023).